OpenStudy (anonymous):

I've worked through this over and over again and still can't figure it out can someone help? Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

\[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x}\]

OpenStudy (anonymous):

and he got \[-10x \sqrt{2x}\]

OpenStudy (anonymous):

I see what he did. If you simplify that 2nd term you get 2xsqrt(2x) 1st term doesn't simplify Last term simplifies to -15sqrt(2x) You can combine the first and last terms, but you have to leave the 2nd term separate because of that x in front of the square root. Nakim ignored the x

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

sure, no problem!

OpenStudy (jdoe0001):

\(\bf 3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x} \implies 3\sqrt{2x}-5\sqrt{18x}+\color{red}{x \sqrt{8x}}\)

OpenStudy (anonymous):

so like bangkokGarrett said, he ignored the x

OpenStudy (jdoe0001):

yes